OpenStudy (anonymous):

i am going yo ask the same question again for help let honestly see if someone will help me. 3x(4 x + x3) simply the equation

OpenStudy (xkehaulanix):

This isn't exactly an equation since there's no equals sign. Is the equation supposed to be \[3x(4x + x^3)\]? You need to multiply what's inside the parenthesis by the 3x on the outside.

OpenStudy (anonymous):

okay so it like 4x3 = 12 right or is it 4x4x4= 64 then 3x3x3= 27 so the answer is 64,27

OpenStudy (xkehaulanix):

The former was about right, actually. Since the 3x is on the outside, you multiply both the 4x and the x^3 by it. Take a look at them separately. 4x * 3x = 4 * x * 3 * x You can only multiply like terms. 4 and 3 are both numbers, and x is a variable that you don't know the value of. 4 * 3 * x * x = 12 * x * x So this part comes out to \[12x^2 \] Can you figure out what the full equation will look like?

OpenStudy (anonymous):

that is what i am trying to do? i dont understand how you done it.

OpenStudy (anonymous):

Alternate forms: 3x^4 + 12x^2 x^2(3x^2+12) 3x^2(x^2+4) Real Root: x = 0 Domain: All real numbers Range: All non-negative real numbers

OpenStudy (xkehaulanix):

Which part did you get confused on? I can try to explain it a little better.

OpenStudy (anonymous):

i know that 3x4=12 12x2 =

OpenStudy (xkehaulanix):

Wait, is it \[3*(4x + x^3)\] or \[3x*(4x+x^3)\]?

OpenStudy (anonymous):

Simplified expression: 21x^2

OpenStudy (anonymous):

3x^4+12x^2